OpenStudy (anonymous):

if f(x) = 5x - 2 then f(-2) = ?

OpenStudy (anonymous):

y=5(-2)-2, evaluate that

OpenStudy (anonymous):

f(-2) = 5(-2) - 2 = -10 - 2 = -12

OpenStudy (anonymous):

thanksss.